OpenStudy (anonymous):

1. (03.03 MC) The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) −1 −11 0 −1 1 9 g(x) g(x) = 5x + 1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has the least y-intercept? Justify your answer. (4 points) (10 points)

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (mathstudent55):

Let's start with part A, the second function, g(x) = 5x + 1 By looking at function g, can you tell what its slope is?

OpenStudy (anonymous):

THE SLOPE IS 5

OpenStudy (mathstudent55):

Correct. Now we need to find the slope of function f, so we can compare slopes. Do you know how to find the slope of a line given 2 points on the line?

OpenStudy (anonymous):

RISE OVER RUN SO DIVIDE THE TWO NUMBERS???

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (mathstudent55):

What two numbers? You are correct. Slope = rise/run. We need to find the rise and the run.

OpenStudy (mathstudent55):

Take any 2 points on the line. Use any two points in the table of function f.

OpenStudy (mathstudent55):

We are given 3 points on function f. Choose any 2. Which two do you want to use?

OpenStudy (anonymous):

the first two

OpenStudy (mathstudent55):

Ok. The rise is the difference in y. Rise means up or down, just like y does. Subtract the y-coordinates of those two points.

OpenStudy (anonymous):

so -11 - (-1)

OpenStudy (mathstudent55):

Here are the first two points. (−1, −11) (0, −1) We find the rise by subtracting the y-coordinates of the two points. Rise = -11 - (-1) = -11 + 1 = -10

OpenStudy (mathstudent55):

Exactly.

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

i will also fan u for your help

OpenStudy (mathstudent55):

The run is the difference in x. Run means right or left, just like x does. Subtract the x-coordinates of those two points. The important thing here is to subtract the x-coordinates in the same order you did the y-coordinates.

OpenStudy (mathstudent55):